Heat of combustion of Ch4, C2H4 and C2H6 are -890,-1411 and -1560 kJ/mol respectively. Which has the lowest calorific value in kJ/g?
  1. Ch4
  2. C2H4
  3. C2H6
  4. All of these

Grade:11

4 Answers

Arun
5 years ago
Dear Abhishek
 
Calorific Value of CH4 = – 890 / 16 =  – 55.53 Kj/ mol
 
Calorific Value of C2H4 =  – 1411 / 28 =  – 50.36 Kj/mol
 
Calorific Value of C2H6 =  – 1560 / 30 =  – 52 Kj / mol
 
Hence C2H4 has least Calorific Value
radhika jaswal
3 years ago
 
Calorific Value of CH4 = – 890 / 16 =  – 55.53 Kj/ mol
 
Calorific Value of C2H4 =  – 1411 / 28 =  – 50.36 Kj/mol
 
Calorific Value of C2H6 =  – 1560 / 30 =  – 52 Kj / mol
 
Hence C2H4 has least Calorific Value
Aarzoo sharma
2 years ago
A value having negative sign has very low value so according to that ch4 should have lower value then why c2h4
Dr Iruwa
2 years ago
My name is Tira  I'm here to make a review about how Dr Mahnlangu helped me in fixing my marriage and also to those out there seeking for help and advice on how to save your broken marriage or relationship, few months ago I was having issues with my husband I noticed he has been having an affair with another young lady in his place of work I confronted him but he got angry and say to my face that it’s true, as if that was not enough he went ahead by telling me he doesn’t love me anymore that he wanted a divorce I was to devastated and confused, I pleaded with him that I don’t want a divorce but he insisted,before then I had heard about Dr Mahlangu while browsing on the internet, how he use to help people with broken marriages and relationships, I gave it a thought and I decided to contact him, he promise to put a smile on my face again, I followed his instructions and after 48 hours I was so surprised to see my husband kneeling before me and apologizing for what he had said about the divorce, this is something I never expected, I thank Dr mahlangu for what he has done for me. Now, I am living happily with my husband and he has stopped seeing the other woman. Please if you are having issues like this I advise you to contact Dr Mahlangu, he can help you as well.
whataspp : +27787390989